OpenStudy (anonymous):

Eliminate the parameter. x = 5 cos t, y = 5 sin t

OpenStudy (anonymous):

\[\cos t=\frac{ x }{ 5 },\sin t=\frac{ y }{ 5 } squareand add\] you will get the answer.

OpenStudy (anonymous):

x^2+y^2=25?

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

Do mind demonstrating to me how you actually got the answer, I want to compare my way with yours, if you do not mind. Also, I want to Thank You for helping me with the previous question you aided me with earlier.

OpenStudy (anonymous):

\[\cos ^{2}t+\sin ^{2}t=\left( \frac{ x }{5 } \right)^{2}+\left( \frac{ y }{ 5 } \right)^{2}\] \[1=\frac{ x ^{2} }{ 25 }+\frac{ y ^{2} }{ 25} ,or x ^{2}+y ^{2}=25\]

OpenStudy (anonymous):

Thank You !